# Simulation of the thermocapillary assembly of a colloidal cluster during the evaporation of a liquid film in an unevenly heated cell

### v2

@@ Abstract
The control of the thermocapillary assembly of colloidal particle clusters is important for a variety of applications, including the creation of photonic crystals for microelectronics and optoelectronics, membrane formation for biotechnology, and surface cleaning for laboratory-on-chip devices. It is important to understand the main mechanisms that influence the formation of such clusters. This article considers a two-dimensional mathematical model describing the transfer of particles by a thermocapillary flow in an unevenly heated cell during the evaporation of a liquid. This gave us the opportunity to study one of the main processes that triggers the formation of a particle cluster. Whether the particle will move with the flow or stop at the heater, becoming the basis for the cluster, is determined by the ratio between gravity and the drag force. The results of numerical calculations show that, for small particle concentrations, their fraction entering the cluster decreases as the volumetric heat flux density $Q$ increases. The reason for this is an increase in the thermocapillary flow with an increase in the volumetric heat flux $Q$. It reduces the probability of particles entering the cluster.
@@ I Introduction
In experiment  Malla, Bhardwaj, and Neild (2019) the effect of particle size and the temperature gradient of the substrate on the effect of coffee rings was studied. In room conditions, an annular deposit formed in place of the dried droplet due to particle motion under the action of capillary flow towards the contact line. The particle size affected the contact line mode (‘‘pinning’’ or ‘‘stick–slip motion’’). The temperature gradient of the substrate led to the appearance of the Marangoni flow, which also affected the transfer of particles. This led to an asymmetry in the shape of the deposit or its different geometry Malla, Bhardwaj, and Neild (2019). An increase in the substrate temperature relative to the ambient temperature led to an increase in the width of the annular deposit (with the presence of a central depleted zone) up to the formation of a continuous spot Lama, Satapathy, and Basavaraj (2020). In another experiment with a droplet, with increasing substrate temperature, a transition from a precipitate in the form of a ‘‘coffee ring’’ to a ‘‘coffee eye’’ was observed. Last one geometry included a narrow annular deposit and a massive central spot inside it Li et al. (2015). The simulation results showed that when the substrate temperature increases, the same change in the geometry of the deposit occurs Zhang et al. (2021). Numerical simulations and experiments were conducted for the case of partial heating of a substrate, where a drop of water containing polystyrene particles was clamped in a Hele-Shaw cell between side plates positioned perpendicular to the horizontal surface of the substrate Thokchom, Majumder, and Singh (2015). This cell shape made it possible to simplify the problem and move on to considering it in 2D. For different heating modes, the flow structure was visualized by the PIV method and calculated using the equations of hydrodynamics. In addition, experiments were conducted on 3D droplets to obtain images of the deposit geometry. Based on these images, particle concentration profiles (the proportion of particles in a local subdomain relative to the total number of particles) were plotted relative to the radial coordinate for different heating modes Thokchom, Majumder, and Singh (2015). The effect of substrate temperature and concentration of aluminum oxide nanoparticles on deposit geometry at site of dried droplet was studied experimentally Liu et al. (2020). At a certain range of values for the key parameters, a formation of deposit in the shape of a double ring was observed. This deposit includes a ‘‘coffee ring’’ and an inner ring Liu et al. (2020). A binary mixture of particles of different sizes was used in an experiment with a drop drying on a heated substrate Parsa et al. (2017). Depending on the temperature of the substrate, five different deposition patterns could be observed: a relatively uniform pattern enclosed by a disk-shaped ring, a nearly non-uniform pattern inside a thick outer ring, a ‘‘dual-ring’’ pattern, a ‘‘rose-like’’ pattern, and a set of concentric rings corresponding to ‘‘stick–slip’’ pattern. The annular deposit in the case of a binary mixture consisted of two sub-rings separated by a small gap. The outer ring consisted of small particles, while the inner ring usually consisted of large particles Parsa et al. (2017). The authors Patil, Bhardwaj, and Sharma (2018) conducted similar experiments and additionally studied the effect of a parameter related to the ratio of the diameter of small particles to that of large ones, $d_{s}/d_{l}$. The resulting phase diagram showed at what values of the parameters $T_{s}$ and $d_{s}$/$d_{l}$ self-sorting of particles by size within the ring occurs Patil, Bhardwaj, and Sharma (2018). For uneven heating of the open cell with liquid, a copper rod was connected to the Peltier element and mounted in the central part of the bottom of the cell Al-Muzaiqer et al. (2021a). The resulting temperature gradient on the free surface of the liquid led to the appearance of a thermocapillary fluid flow. The transfer of particles by the flow influenced the creation of a cluster of particles in the central region of the cell. The numerically calculated flow velocity in order of magnitude was in good agreement with the measurement results obtained by the PIV method Al-Muzaiqer et al. (2021a). The proposed 1D model allows to consider only the horizontal flow velocity averaged over the height of the liquid film. But it doesn’t allow one to determine the 2D structure of the flow Al-Muzaiqer et al. (2021a). The simulation showed that, over time, the concentration of particles increases in the area of the heater. This technology can be used to control microparticles, including solids, polymers, cells, micelles, proteins, and microemulsions that are suspended in a liquid medium (see Refs. in paper Al-Muzaiqer et al. (2021b)). The results of numerical calculations of the flow of pure liquid (without particles) in an evaporating droplet predicted the possibility of controlling the flow structure by changing the point of local heating of the substrate Lee, Choi, and Lee (2022). Another numerical study Jain, Denner, and van Wachem (2024) showed that at temperature $T_{s}=$ 25C capillary flow prevails in the evaporating droplet. The Marangoni heat flow dominates at temperature $T_{s}=$ 50C. In addition, the effect of the ratio of particle density to liquid density, $\rho_{p}/\rho_{l}$, on particle motion in a fluid flow was studied Jain, Denner, and van Wachem (2024). The effect of the nanoparticle concentration on the structure of deposits remaining after droplet drying on a heated substrate ($T_{s}=$ 80C) was studied both numerically and experimentally Kumar, Sajith, and Sathian (2021). At a relatively low concentration of the solution, an entire spot was formed. When the concentration of nanoparticles was relatively high, the coffee ring effect was observed Kumar, Sajith, and Sathian (2021). In Ref. Goncharova and Bekezhanova (2024), the effect of two heaters mounted in a substrate on the structure of the flow of pure liquid in an evaporating film was studied using the proposed mathematical model. The corresponding flow structure was shown for different heating modes. The effect of particle sedimentation on the sediment structure after drying of the suspended droplet on substrates of different wettability was investigated in Ref. Kumar, Thampi, and Basavaraj (2020). The obtained precipitation structures were compared with the structures for the case of a sessile drop. By adjusting the ambient temperature, it is possible to control the formation of an array of microneedles when the polymer solution film dries Mansoor and Stoeber (2011). Such devices are useful for transdermal drug delivery system in medicine.
@@ II.1 Physical problem statement
Let us consider a liquid located in an open cylindrical cell, in the lower part of which a heater is mounted (Fig. 1). The cell consists of a substrate of plexiglass, with a photopolymer ring glued on top of it. A copper rod is mounted in the center of the bottom of the cell, which is connected to the heated side of a Peltier element. As working materials, polystyrene microspheres and volatile isopropyl alcohol (isopropanol) were used. The problem has axial symmetry and is solved in a cylindrical coordinate system $(r,\,z)$. The geometrical parameters and properties of materials are summarized in Table 1 and  2, their values are taken from Ref. Al-Muzaiqer et al. (2021a)
@@
$R_{in}$ Inner radius of the cell 10
@@
$L$ Isopropanol heat of vaporization $75\times 10^{4}$ J///kg
$M$ Isopropanol molar mass 0.06 kg///mol
$\rho_{l}$ Isopropanol density 786 kg///m3
$\rho_{s}$ Plexiglass density 2500 kg///m3
$\rho_{p}$ Polystyrene particle density 1050 kg///m3
$\rho_{\mathrm{rod}}$ Copper density 8960 kg///m3
@@
$c_{l}$ Isopropanol specific heat capacity 2605 J///(kg K)
$c_{s}$ Plexiglass specific heat capacity 780 J///(kg K)
$c_{\mathrm{rod}}$ Copper specific heat capacity 385 J///(kg K)
$\kappa_{l}$ Isopropanol thermal conductivity 0.13 W///(m K)
$\kappa_{s}$ Plexiglass thermal conductivity 0.748 W///(m K)
$\kappa_{\mathrm{rod}}$ Copper thermal conductivity 400 W///(m K)
$\alpha_{\mathrm{air}}$ Air heat transfer coefficient 5.6 W///(m2 K)
$\alpha_{l}$ Isopropanol heat transfer coefficient 216 W///(m2 K)
$\alpha_{s}$ Plexiglass heat transfer coefficient 2.5 W///(m2 K)
$\sigma$ Surface tension $22\times 10^{-3}$ N///m
$\sigma^{\prime}$ Isopropanol surface tension temperature coefficient, $\sigma^{\prime}=\partial\sigma/\partial T$ $-8\times 10^{-5}$ N///(m K)
@@ II.2 Mathematical model
where $Q$ is the volumetric heat flux density (W/m3).
@@
where the drag force is $\vec{F}_{d}=m_{p}\left(\vec{v}-\vec{u}\right)/\tau_{p}$, particle response time is $\tau_{p}=\rho_{p}d_{p}^{2}/(18\mu)$, particle diameter is $d_{p}=2r_{p}$, the gravity force is $\vec{F}_{g}=m_{p}\vec{g}$ ($\vec{g}$ is the acceleration due to gravity), the buoyant force (Archimedes’ principle) is $\vec{F}_{A}=-\rho_{l}V_{p}\vec{g}$, and particle volume is $V_{p}=4\pi r_{p}^{3}/3$.
@@ II.3 Initial and boundary conditions
At the boundary ‘‘liquid–cell’’, the fluid flow velocity is zero, $\vec{v}=0$, due to the presence of viscous friction and no penetration through the substrate and cell wall. The following boundary conditions are used on a free surface Barash (2009),
@@
where $\vec{\tau}$ is the tangential to the surface, $P_{L}=\sigma\left(1/R_{1}+1/R_{2}\right)$ is the Laplace pressure, $R_{1},\,R_{2}$ are the main radii of curvature on the free surface, $d\sigma/d\tau=\sigma^{\prime}\,\partial T/\partial\tau$, $\phi$ is the angle between the vector $\vec{\tau}$ and $r$-axis (or angle between the vector $\vec{n}$ and $z$-axis) Turchaninova et al. (2021). Calculation details of $\partial\phi/\partial\tau$ are described in Ref. Barash et al. (2009). In addition, on the axis of symmetry $v_{r}=0$ and $\partial v_{z}/\partial r=0$ Turchaninova et al. (2021).
@@ III Results and discussion
A series of computational experiments have been carried out for different values of volumetric heat flux, $Q$, and the number of particles in the system, $N_{\mathrm{tot}}$. Each experiment was repeated five times, after which the result was averaged over all trials. The standard deviation was used to calculate the error. The calculation time depends on $N_{\mathrm{tot}}$. It varies from a few hours to one or two days (Intel(R) Core(TM) i7-5820K, CPU 3.3 GHz, RAM 32 Gb, SSD 128 Gb (system), SSD 1 Tb (data)).
@@
Figure 3: Dependence of the maximum thermocapillary flow velocity on time for several values $Q$.
@@
The thermocapillary effect causes the free surface of the liquid to curve due to uneven heating (Fig. 4). This is due to a surface tension gradient caused by a temperature difference. By the cell wall, the liquid layer becomes higher as the liquid flows there, because the surface tension is higher there due to a lower temperature (Fig. 5). In the central area of the cell, the thickness of the liquid layer decreases, which leads to thermocapillary rupture and the formation of a dry spot (Fig. 5). In the area of the heater, there is an additional decrease in the thickness of the liquid layer due to uneven evaporation. This is because the higher the temperature of the liquid, the more actively it evaporates (Fig. 6). Over time, as the liquid evaporates, its height decreases (Fig. 4).
@@
Figure 8: Particle dynamics at $Q=$ 84 MW/m3 and $N=125$ (multimedia available online).

### v3

@@ Abstract
The control of the thermocapillary assembly of colloidal particle clusters is important for a variety of applications, including the creation of photonic crystals for microelectronics and optoelectronics, membrane formation for biotechnology, and surface cleaning for laboratory-on-chip devices. The aim of the current work is to understand the main mechanisms influencing the assembly of such clusters, which are the object of our research. Here we consider a two-dimensional mathematical model describing the transfer of particles by a thermocapillary flow in an unevenly heated cell during the evaporation of a liquid. This gave us the opportunity to study one of the main processes that triggers the formation of a particle cluster. Whether the particle will move with the flow or stop at the heater, becoming the basis for the cluster, is determined by the ratio between gravity and the drag force. The results of numerical calculations show that, for small particle concentrations, their fraction entering the cluster decreases as the volumetric heat flux density $Q$ increases. The reason for this is an increase in the thermocapillary flow with an increase in the volumetric heat flux $Q$. In our conclusion, it reduces the probability of particles entering the cluster.
colloidal cluster, particles assembly, thermocapillary flow, uneven heating, liquid evaporation in open cell
@@ I Introduction
In experiment Malla, Bhardwaj, and Neild (2019) the effect of particle size and the temperature gradient of the substrate on the effect of coffee rings was studied. In room conditions, an annular deposit formed in place of the dried droplet due to particle motion under the action of capillary flow towards the contact line. The particle size affected the contact line mode (‘‘pinning’’ or ‘‘stick–slip motion’’). The temperature gradient of the substrate led to the appearance of the Marangoni flow, which also affected the transfer of particles. This led to an asymmetry in the shape of the deposit or its different geometry Malla, Bhardwaj, and Neild (2019). An increase in the substrate temperature relative to the ambient temperature led to an increase in the width of the annular deposit (with the presence of a central depleted zone) up to the formation of a continuous spot Lama, Satapathy, and Basavaraj (2020). In another experiment with a droplet, with increasing substrate temperature, a transition from a precipitate in the form of a ‘‘coffee ring’’ to a ‘‘coffee eye’’ was observed. Last one geometry included a narrow annular deposit and a massive central spot inside it Li et al. (2015). The simulation results showed that when the substrate temperature increases, the same change in the geometry of the deposit occurs Zhang et al. (2021). Numerical simulations and experiments were conducted for the case of partial heating of a substrate, where a drop of water containing polystyrene particles was clamped in a Hele-Shaw cell between side plates positioned perpendicular to the horizontal surface of the substrate Thokchom, Majumder, and Singh (2015). This cell shape made it possible to simplify the problem and move on to considering it in 2D. For different heating modes, the flow structure was visualized by the PIV method and calculated using the equations of hydrodynamics. In addition, experiments were conducted on 3D droplets to obtain images of the deposit geometry. Based on these images, particle concentration profiles (the proportion of particles in a local subdomain relative to the total number of particles) were plotted relative to the radial coordinate for different heating modes Thokchom, Majumder, and Singh (2015). The effect of substrate temperature and concentration of aluminum oxide nanoparticles on deposit geometry at site of dried droplet was studied experimentally Liu et al. (2020). At a certain range of values for the key parameters, a formation of deposit in the shape of a double ring was observed. This deposit includes a ‘‘coffee ring’’ and an inner ring Liu et al. (2020). A binary mixture of particles of different sizes was used in an experiment with a drop drying on a heated substrate Parsa et al. (2017). Depending on the temperature of the substrate, five different deposition patterns could be observed: a relatively uniform pattern enclosed by a disk-shaped ring, a nearly non-uniform pattern inside a thick outer ring, a ‘‘dual-ring’’ pattern, a ‘‘rose-like’’ pattern, and a set of concentric rings corresponding to ‘‘stick–slip’’ pattern. The annular deposit in the case of a binary mixture consisted of two sub-rings separated by a small gap. The outer ring consisted of small particles, while the inner ring usually consisted of large particles Parsa et al. (2017). The authors Patil, Bhardwaj, and Sharma (2018) conducted similar experiments and additionally studied the effect of a parameter related to the ratio of the diameter of small particles to that of large ones, $d_{s}/d_{l}$. The resulting phase diagram showed at what values of the parameters $T_{s}$ and $d_{s}$$d_{l}$ self-sorting of particles by size within the ring occurs Patil, Bhardwaj, and Sharma (2018). For uneven heating of the open cell with liquid, a copper rod was connected to the Peltier element and mounted in the central part of the bottom of the cell Al-Muzaiqer et al. (2021a). The resulting temperature gradient on the free surface of the liquid led to the appearance of a thermocapillary fluid flow. The transfer of particles by the flow influenced the creation of a cluster of particles in the central region of the cell. The numerically calculated flow velocity in order of magnitude was in good agreement with the measurement results obtained by the PIV method Al-Muzaiqer et al. (2021a). The proposed 1D model allows to consider only the horizontal flow velocity averaged over the height of the liquid film. But it doesn’t allow one to determine the 2D structure of the flow Al-Muzaiqer et al. (2021a). The simulation showed that, over time, the concentration of particles increases in the area of the heater. This technology can be used to control microparticles, including solids, polymers, cells, micelles, proteins, and microemulsions that are suspended in a liquid medium (see Refs. in paper Al-Muzaiqer et al. (2021b)). The results of numerical calculations of the flow of pure liquid (without particles) in an evaporating droplet predicted the possibility of controlling the flow structure by changing the point of local heating of the substrate Lee, Choi, and Lee (2022). Another numerical study Jain, Denner, and van Wachem (2024) showed that at temperature $T_{s}=$ 25C capillary flow prevails in the evaporating droplet. The thermal Marangoni flow dominates at temperature $T_{s}=$ 50C. In addition, the effect of the ratio of particle density to liquid density, $\rho_{p}/\rho_{l}$, on particle motion in a fluid flow was studied Jain, Denner, and van Wachem (2024). The effect of the nanoparticle concentration on the structure of deposits remaining after droplet drying on a heated substrate ($T_{s}=$ 80C) was studied both numerically and experimentally Kumar, Sajith, and Sathian (2021). At a relatively low concentration of the solution, an entire spot was formed. When the concentration of nanoparticles was relatively high, the coffee ring effect was observed Kumar, Sajith, and Sathian (2021). In Ref. Goncharova and Bekezhanova (2024), the effect of two heaters mounted in a substrate on the structure of the flow of pure liquid in an evaporating film was studied using the proposed mathematical model. The corresponding flow structure was shown for different heating modes. The effect of particle sedimentation on the precipitation structure after drying of the pendant droplet on substrates of different wettability was investigated in Ref. Kumar, Thampi, and Basavaraj (2020). The obtained precipitation structures were compared with the structures for the case of a sessile drop. By adjusting the ambient temperature, it is possible to control the formation of an array of microneedles when the polymer solution film dries Mansoor and Stoeber (2011). Such devices are useful for transdermal drug delivery system in medicine.
@@ II.1 Physical problem statement
Let us consider a liquid located in an open cylindrical cell, in the lower part of which a heater is mounted (Fig. 1). The cell consists of a substrate of plexiglass, with a photopolymer ring glued on top of it. A copper rod is mounted in the center of the bottom of the cell, which is connected to the heated side of a Peltier element. As working materials, polystyrene microspheres and volatile isopropyl alcohol (isopropanol) were used. The problem has axial symmetry and is solved in a cylindrical coordinate system (r,z)(r,\,z). The geometrical parameters and properties of materials are summarized in Table 1 and 2, their values are taken from Ref. Al-Muzaiqer et al. (2021a)
@@
$R_{\mathrm{in}}$ Inner radius of the cell 10
@@
$L$ Isopropanol heat of vaporization $75\times 10^{4}$ J//kg
$M$ Isopropanol molar mass 0.06 kg//mol
$\rho_{l}$ Isopropanol density 786 kg//m3
$\rho_{s}$ Plexiglass density 2500 kg//m3
$\rho_{p}$ Polystyrene particle density 1050 kg//m3
$\rho_{\mathrm{rod}}$ Copper density 8960 kg//m3
@@
$c_{l}$ Isopropanol specific heat capacity 2605 J//(kg K)
$c_{s}$ Plexiglass specific heat capacity 780 J//(kg K)
$c_{\mathrm{rod}}$ Copper specific heat capacity 385 J//(kg K)
$\kappa_{l}$ Isopropanol thermal conductivity 0.13 W//(m K)
$\kappa_{s}$ Plexiglass thermal conductivity 0.748 W//(m K)
$\kappa_{\mathrm{rod}}$ Copper thermal conductivity 400 W//(m K)
$\alpha_{\mathrm{air}}$ Air heat transfer coefficient 5.6 W//(m2 K)
$\alpha_{l}$ Isopropanol heat transfer coefficient 216 W//(m2 K)
$\alpha_{s}$ Plexiglass heat transfer coefficient 2.5 W//(m2 K)
$\sigma_{0}$ Surface tension $22\times 10^{-3}$ N//m
$\sigma^{\prime}$ Isopropanol surface tension temperature coefficient, $\sigma^{\prime}=\partial\sigma/\partial T$ $8\times 10^{-5}$ N//(m K)
@@ II.2 Mathematical model
where $Q$ is the volumetric heat flux density (W/m3). This model uses a single temperature notation $T$ for each subdomain, since Eqs. (3)–(5) can be rewritten as a single general equation, where the parameter values change depending on the subdomain for which it is being calculated. Furthermore, the convective term contained in Eq. (3) and the source term $Q$ from Eq. (5) can automatically be set to zero in those subdomains where they are not required.
@@
where the drag force is $\vec{F}_{d}=m_{p}\left(\vec{v}-\vec{u}\right)/\tau_{p}$, particle response time is $\tau_{p}=\rho_{p}d_{p}^{2}/(18\mu)$, particle diameter is $d_{p}=2r_{p}$, the gravity force is $\vec{F}_{g}=m_{p}\vec{g}$ ($\vec{g}$ is the acceleration due to gravity), the buoyant force (Archimedes’ principle) is $\vec{F}_{A}=-\rho_{l}V_{p}\vec{g}$, and particle volume is $V_{p}=4\pi r_{p}^{3}/3$. In Eq. (6), we neglect the added mass, since the Stokes number $\mathrm{St}=\tau_{p}/\tau_{f}\approx 3\cdot 10^{-4}\ll 1$. Here, the characteristic flow time is calculated by the formula $\tau_{f}=\rho_{p}h_{0}^{2}/\mu\approx$ 0.2 s.
@@ II.3 Initial and boundary conditions
Here, the subscripts for $T$ are used solely for ease of perception and understanding of the physical meaning of the boundary conditions. It should be remembered that there is only one temperature function in the model.
@@
At the boundary ‘‘liquid–cell’’, the fluid flow velocity is zero, $\vec{v}=0$, due to the presence of viscous friction and no penetration through the substrate and cell wall. The following boundary conditions are used on a free surface of the liquid Barash (2009),
@@
where $\vec{\tau}$ is a vector tangent to the surface, directed perpendicularly to the right when looking along the normal vector direction (the arrow above the directional derivative is omitted for brevity), $P_{L}=\sigma\left(1/R_{1}+1/R_{2}\right)$ is the Laplace pressure, $R_{1},\,R_{2}$ are the main radii of curvature on the free surface, $d\sigma/d\tau=\sigma^{\prime}\,\partial T/\partial\tau$, $\phi$ is the angle between the vector $\vec{\tau}$ and $r$-axis (or angle between the vector $\vec{n}$ and $z$-axis) Turchaninova et al. (2021). Calculation details of $\partial\phi/\partial\tau$ are described in Ref. Barash et al. (2009). In addition, on the axis of symmetry $v_{r}=0$ and $\partial v_{z}/\partial r=0$ Turchaninova et al. (2021).
@@
As the contact line is well away from the centre where the cluster is assembled in the cell, we use a simple boundary condition, $\theta=\pi/2$, for the contact angle.
@@ II.4 Closing relations
In this model the dependence of surface tension on temperature is described by the linear approximation $\sigma=\sigma_{0}+\sigma^{\prime}(T_{l}-T_{0})$.
@@ III Results and discussion
A series of computational experiments have been carried out for different values of volumetric heat flux, $Q$, and the number of particles in the system, $N_{\mathrm{tot}}$. Each simulation of particle motion was repeated five times, after which the result was averaged over all trials. The standard deviation was used to calculate the error (the initial position of the particles was randomly generated), see Fig. 2. The temperature and flow velocity fields were calculated once for each considered value of the parameters $N_{\mathrm{tot}}$ and $Q$. The calculation time depends on $N_{\mathrm{tot}}$. It varies from a few hours to one or two days (Intel(R) Core(TM) i7-5820K, CPU 3.3 GHz, RAM 32 Gb, SSD 128 Gb (system), SSD 1 Tb (data)).
@@
Figure 3: Dependence of the maximum thermocapillary flow velocity on time for several values $Q$ (the numerical calculation error is less than the marker size).
@@
The thermocapillary effect causes the free surface of the liquid to curve due to uneven heating (Fig. 4). It is in qualitative agreement with the results from the 1D model Al-Muzaiqer et al. (2021a). This is due to a surface tension gradient caused by a temperature difference. By the cell wall, the liquid layer becomes higher as the liquid flows there, because the surface tension is higher there due to a lower temperature (Fig. 5). In the central area of the cell, the thickness of the liquid layer decreases, which leads to thermocapillary rupture and the formation of a dry spot (Fig. 5). In the area of the heater, there is an additional decrease in the thickness of the liquid layer due to uneven evaporation. This is because the higher the temperature of the liquid, the more actively it evaporates (Fig. 6). Over time, as the liquid evaporates, its height decreases (Fig. 4). At $Q=$ 84 MW/m3, the temperature difference $\Delta T$ is approximately 34 K. Let us determine the value of the Marangoni number, $\mathrm{Ma}=-\sigma^{\prime}\Delta T\,h_{0}/(\mu D_{T})\approx 1.258\times 10^{4}$, where the thermal diffusivity coefficient is $D_{T}=\kappa_{l}/(\rho_{l}c_{l})\approx 6.3\times 10^{-8}$ m2/s. At such relatively large temperature differences considered here, strong thermocapillary convection arises, which is described by the proposed model. However, it should be noted that the temperature must not exceed a critical value at which liquid boiling will occur. Boiling can have a negative impact on particle self-assembly. Furthermore, the model does not describe such a process.
@@
Figure 8: Particle dynamics at $Q=$ 84 MW/m3 and $N=125$ (multimedia available in Supplementary Material).
To understand which of the two processes, evaporation or thermocapillary flow, has a greater influence on the thinning of the liquid layer near the heater, let us compare two characteristic times. We estimate the characteristic evaporation time as $t_{e}=h_{0}/v_{e}$, where the characteristic evaporation velocity is $v_{e}=J_{\mathrm{max}}/\rho_{l}$ (for $Q=$ 84 MW/m3, $J_{\mathrm{max}}\approx 3.25\times 10^{-3}$ kg/(m2s) according to Fig. 6). The characteristic time of thermocapillary convection can be estimated by the formula $t_{M}=h_{0}/v_{M}$, where the characteristic Marangoni flow velocity is $v_{M}\approx 2.5\times 10^{-2}$ m/s (Fig. 7b). Substituting the parameter values, we obtain $t_{e}\approx 169$ s and $t_{M}\approx 0.03$ s. The rapid thinning of the liquid film in the central part of the cell is associated with thermocapillary convection, since $t_{M}\ll t_{e}$. A dry spot forms as a result of a thermocapillary rupture, where the central cluster ends up outside the liquid. Subsequent evaporation of the remaining liquid from the cell and the transport of particles suspended in it do not affect the size of this cluster. Let us also estimate the thermal relaxation time, $t_{r}=h_{0}^{2}/D_{T}\approx$ 7.8 s. From the relation $t_{M}\ll t_{r}$, it follows that heat transfer in the liquid due to convection dominates over thermal conduction.